A sailboat takes 4 hours to travel 128 km going upstream. The return trip takes 2 hours going down stream. What is the rate of the boat in still water and what is the rate of the current?
Answer by TimothyLamb(4379)

x = speed of boat in still water
y = speed of current
---
s = d/t
---
upstream:
x - y = 128/4
---
downstream:
x + y = 128/2
---
put the system of linear equations into standard form
---
x - y = 32
x + y = 64
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = speed of boat in still water = 48 kph
y = speed of current = 16 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php